OpenStudy (anonymous):

explain how to completely factor 2x^2 − 8x − 42.

myininaya (myininaya):

find two factors of 2(-42) that have product 2(-42) and have sum -8; 2(-42)=4(-21)=12(-7)=6(-14) -8=-14+6 -8x=-14x+6x so replace -8x with -14x+6x and then factor by grouping

OpenStudy (anonymous):

2x^2 - 8x - 42 --> 2(x^2 - 4x - 21) --> 2 (x +3)(x -7) --> x = -3 or x = 7

OpenStudy (anonymous):

Factor out a 2: 2(x²-4x-21) Product-sum of -21&-4 is -7 and +3, so it becomes: 2(x-7)(x+3)

OpenStudy (anonymous):

2(x^2-4x-21)=2(x^2-7x+3x-21)=2(x(x-7)+3(x-7))=2(x-7)(x+3)

myininaya (myininaya):

lol i guess you factor out a 2 first

myininaya (myininaya):

gj you guys

OpenStudy (anonymous):

haha thanks myininaya!